# HateDebias: On the Diversity and Variability of Hate Speech Debiasing

### v1

@@ 3.2 Debiasing Performance Forgetting of Continuous Debiasing
Initially, we analyze the results of existing methods to determine whether the continuous debiasing of newly emerging attributes will affect the overall fairness of the model, that is, whether the debiasing performance of the model suffers from the catastrophic forgetting problem in the continuous debiasing task. As shown in Table 5, comparing five MTL-based methods (static fairness methods) with five Fine-tune-based methods (continual debiasing methods), it can be found that continual debiasing tasks are prone to be more biased than static debiasing tasks. Fine-tune-based methods can be regarded as static debiasing methods adapted to continuous scenes. These methods perform poorly in continuous scenarios, indicating models struggle with ever-changing biases. The HateDebias dataset closely mirrors real-world bias occurrences, validating that static methods are inadequate under multiple, continual biases. This suggests that in practical applications, more adaptive debiasing methods are needed to improve the overall performance and fairness of models. This underscores the rationale and necessity of the HateDebias dataset proposed in this paper. It closely resembles the occurrence of bias in real-world scenarios, simulating multiple, continuously changing biases, and provides a real environment closer to practical applications.
3.3 Performance of Different Methods on HateDebias
@@
Our experimental findings, detailed in Table 5, indicate that our approach yields substantial improvements in continual debiasing tasks. Static debiasing methods MTL, which can be considered as an upper-bound followed by [25, 29], show significant debiasing effects when dealing with multiple sensitive attributes. In contrast, fine-tuning-based continuous debiasing methods have limited debiasing capabilities, indicating their ineffectiveness in addressing continuous bias. Additionally, the continuous debiasing effect using the ER, REPINA and IRD methods is worse, indicating an inability to effectively reduce bias. These methods are ineffective because the HateDebias dataset closely mirrors real-world bias occurrences, simulating multiple and continuously changing biases. Experimental results show that existing methods perform poorly in these complex, dynamic environments. Finally, the experimental results show that our proposed framework significantly reduces bias in continuous debiasing tasks. Especially with supervised debiasing strategies, the model’s bias is notably reduced at the cost of slight performance trade-offs in downstream tasks. Compared to methods like REPINA, IRD, and TKD, the CLF method demonstrates better debiasing effects and model performance across various debiasing strategies. Experimental results show the proposed method CLF excels in handling multiple and continuously changing biases in complex real-world scenarios, effectively debiasing while maintaining high performance and fairness.
3.4 Performance of Debiasing Preservation Capability
Table 6: The results of debiasing preservation capability.
@@ 3.4 Performance of Debiasing Preservation Capability
$BC=\frac{1}{n-1}\sum_{i=1}^{n-1}FPR_{i}^{n}-FPR_{i}^{i},$ (5)
where $FPR_{i}^{i}$ is the false positive rate of the $i$-th bias in the $i$-th stage. Table 6 presents the results of the debiasing preservation capability. These results demonstrate that our proposed CLF method significantly outperforms others in preserving debiasing effects over time. Specifically, the CLF method achieves a BC value of -1.60, indicating substantial bias reduction and showcasing its effectiveness in mitigating bias. In comparison, the REPINA method shows no improvement, while methods like TKD and IDBR exhibit moderate improvements. The poor performance of the comparative methods indirectly reflects the reasonableness of our dataset in terms of capturing the diversity and variability of real-world biases.
Table 7: Experimental results of different sequence lengths.
@@
3.5 Effects of Different Length Settings on Hatedebias
In practical applications, biases come in various types and forms. To gradually make our dataset approximate real-world scenarios, we experiment with different sequence lengths to evaluate the debiasing effects under different length settings. We compare continuous debiasing tasks with sequence lengths of 2, 3 and 4 to analyze how the diversity and variability of the dataset affect model performance and fairness. As shown in Table  7, we choose CL as the debiasing method and the results reveal several key insights: (1) As the sequence length increases, the performance metrics generally show a decreasing trend. This suggests that longer sequences may introduce more complexity, making it challenging for models to maintain high performance. (2) The AAB values decrease with longer sequence lengths, indicating improved fairness. This trend demonstrates that datasets with longer sequences better capture the diversity and variability of real-world biases, helping models to generalize and debias more effectively. (3) Our proposed method CLF exhibits the best continuous debiasing capability across different sequence lengths, indicating that it can effectively cover datasets in various real-world scenarios.
4 Conclusion
In this paper, we focus on the continuous debiasing task, addressing the problem of debiasing in hate speech scenarios with multi-type and continuously varying bias types. We have introduced a dataset named HateDebiase to benchmark continuous debiasing tasks. The HateDebiase dataset contains 23,276 samples with four bias attributes, corresponding to four sub-datasets within HateDebiase. Based on HateDebiase, we have also proposed a continuous-learning-based framework to eliminate emerging biases while preserving model performance in downstream tasks. We evaluate the debiasing performance of some widely used methods on HateDebiase and demonstrate that the continuous-learning-based method is a promising direction for improving continuous debiasing tasks.
Supplement to
“HateDebias: On the Diversity and Variability of Hate Speech Debiasing”
Appendix organization:
[…]
1
Appendix A1 Related Work
Datasets of Hata Speech Detection.
Static Debiasing.
In the field of Natural Language Processing (NLP), bias mitigation strategies [31, 32, 33] fall into two primary categories. The first involves training models to reduce the influence of sensitive attributes during fine-tuning [34, 35], using techniques like projection-based [36], adversarial [17], or contrastive objectives [37]. This approach requires data specifically annotated for the sensitive attribute, with the debiasing effectiveness dependent on the external corpus quality and the potential introduction of new biases. The second category, task-agnostic training, uses general corpora to address biases, such as by identifying and removing gender biases from encoded representations [38] or enhancing model fairness through selective data exclusion [39], increased dropout rates [40], or equalizing objectives [41]. Additionally, methods like the introduction of fairness-triggering perturbations [42] and equi-tuning for group equivariance in immutable models [43] have been explored, focusing on minimizing differences between standard and debiased model features.
Continuous Debiasing.
Continual Learning.
Continual learning involves updating models with new information while retaining previous knowledge, avoiding catastrophic forgetting [46]. This process entails maintaining generic internal representations for reuse across tasks and dynamically adjusting task-specific parts. Effective strategies include memory-based methods that use prior task exemplars [47, 24] or synthetic data [48], and regularization techniques that minimize changes in learned representations under a fixed architecture [49, 50]. For example, Li and Hoiem [51] proposed "learning without forgetting" (LwF), a technique that introduces regularization to align the model’s output on the current data with that of the previously trained model. Knowledge distillation [52, 53], gradient-based [54] and expansion-based approaches [55] also mitigate forgetting, each targeting the preservation and enhancement of model adaptability over time. Furthermore, some research endeavors have shifted attention toward continual-learning-based methodologies as potential solutions to address model fairness concerns. Churamani et al. [56] introduced a pioneering application of domain-incremental learning as an effective approach for mitigating bias, specifically in the context of enhancing the fairness of facial expression recognition systems. However, there is no work focusing on the continuous debiasing task.
Figure A1: An illustration of stastic debiasing task and continuous debiasing task. In the static debiasing task, the emergence of a new bias requires a complete re-debiasing of all existing biases, resulting in an exponential increase in both training time and costs. Conversely, in the continuous debiasing task, the emergence of a new bias requires only the fine-tuning of this specific bias, eliminating the need for re-training.
Appendix A2 Illustration of Continuous Debiasing and Static Debiasing
The static debiasing primarily focused on adjusting specific biases, as illustrated in Fig. A1. Static debiasing requires re-debiasing all biases whenever new ones appear, consuming considerable training time and resources. Such a cost makes its application impractical in real-world environments where new biases emerge continuously. For instance, if a model is initially trained to mitigate racial bias, and then gender bias becomes apparent, the entire model needs to be retrained to address both biases simultaneously. This retraining process can be highly inefficient and resource-intensive, especially as the number and types of biases continue to grow.
In fact, models in ever-changing environments are expected to adjust targetedly and dynamically to correct these continuously emerging biases, which we define as continuous debiasing, as shown at the bottom of Fig. A1. Continuous debiasing allows for incremental updates to the model to address new biases as they arise without requiring a complete overhaul of the existing bias mitigation strategies. This type of task is significantly more efficient and scalable, making it more suitable for real-world applications where the social and cultural context is constantly evolving.
The importance and rationale of dynamic debiasing lie in its ability to maintain the performance and fairness of models over time. As new biases emerge, they can quickly degrade the effectiveness of a model if not addressed promptly. For example, a social media platform using a hate speech detection model might initially focus on racial slurs. Over time, if there is a rise in hate speech targeting a specific gender or sexual orientation, the model’s effectiveness could diminish unless these new biases are dynamically mitigated. In a continuously changing environment, static debiasing methods fall short as they do not adapt to the new forms of biases that emerge. Continuous debiasing, however, ensures that the models are regularly updated, thus upholding the standards of fairness and reducing the risk of perpetuating new biases. Moreover, dynamic debiasing helps preserve the investment in existing models by extending their useful life and maintaining their relevance without the need for complete retraining.
Appendix A3 Pseudocode of Our Proposed Method
The pseudocode of the implementation process of our proposed continual-learning-based debiasing framework is shown in Algorithm 1.
Algorithm 1 CLF
2:Trained language model encoder $LM^{n}$
3:Load pre-trained language model $LM^{0}$
4:for $i=1,...,n$ do
7:     if $i\geq 2$ then
8:         Calculate $\mathcal{L}_{g}$ and $\mathcal{L}_{s}$ with $LM^{i}$ (Eq. 7-8) $\triangleright$ Bias Information Regularization
10:         $M_{t}$ = []
11:         for $r=1,...,i$ do $\triangleright$ Memory Replay
12:              $M_{t}\leftarrow\gamma|D_{r}|$ (Eq. 11)
13:              $M_{t}\leftarrow M_{t}\cup M_{t}^{r}$
14:         end for
15:     end if
17:     Update model parameters with $\mathcal{L}$
18:end for
Appendix A4 Continual-learning-based Framework
A4.1 Overview
We present a continual-learning-based framework for continual debiasing tasks, aiming to enhance the model’s performance and fairness by promoting the learning of attribute-irrelevant features to effectively mitigate bias. For a given task, our proposed method is adopted to the two-stage training backbone, in terms of the initial training stage and the memory replay stage, where the replay-based method can effectively alleviate the classifier bias. During the initial training stage, the model performs training solely on the current task data, allowing the learning of new task-related knowledge. Subsequently, in the memory replay stage, a subset consisting of representative samples from the previous task data is selected to update the memory. The model is further retrained using the balanced memory bank, facilitating the retrieval and consolidation of previously acquired knowledge. Specifically, after training on $(t-1)$-th $(t\geq 2)$ task, $\gamma|D_{t-1}|$ examples are sampled from the training set $D_{t-1}$ into the memory buffer $M$, where $0\leq\gamma\leq 1$ is the store ratio. Data from $M$ is then merged with the next training set $D_{t}$ when conducting the learning of task $t$. What’s more, in the process of continual learning, an important aspect arises from the presence of shared knowledge across different debiasing tasks. Hence, the model must constantly adapt to the specific knowledge related to each debiasing task while avoiding forgetting the old knowledge throughout the learning process. To tackle this challenge, we disentangle sensitive attribute information from raw text representation to obtain a generic space and bias-specific space, introducing an bias information regularization technique to allow interaction between two spaces. The regularization approach aims to preserve shared knowledge while enabling the model to effectively incorporate bias-specific knowledge for different debiasing tasks.
A4.2 Bias Information Regularization
The operation of disentangled enables the separation of generic and bias-specific information in the hidden representations, facilitating a more refined analysis and understanding of the sentence’s underlying structure and context. The disentangled concept we define allows the model to learn generic information about downstream tasks and specific information about sensitive attributes, respectively. For a given sentence $x$, a pre-trained language model $B(\cdot)$ is employed to derive the hidden representations $h$, encompassing both generic and bias-specific information. Subsequently, we design two disentangled networks, represented by $G(\cdot)$ and $S(\cdot)$, which play a vital role in extracting the generic representation $g$ and sensitive bias-specific representation $s$ from the previously obtained hidden representations $h$. $G(\cdot)$ and $S(\cdot)$ refer to two distinct feed-forward networks. The dimensions of the representations produced by $G(\cdot)$ and $S(\cdot)$ are identical to the hidden representations $h$. This systematic process enables the generic and bias-specific information to disentangle, facilitating a more refined analysis and understanding of the sentence’s underlying structure and context. For the dynamic debiasing task of sensitive attributes, we train a classifier by exploiting information from two representations and allowing different representation spaces to be changed to different extents, maintaining the ability of the models when debiasing the sensitive attributes.
Generic Representation
Bias-specific Representation
$\mathcal{L}_{task}=E_{(x,z)\in D_{t}\cup M_{t}}\mathcal{L}_{CE}(f_{task}(S(B(x)),z),$ (6)
where $z$ is the corresponding debiasing task id for $x$, $L_{CE}(\cdot)$ is the cross entropy loss function and $B(\cdot)$ is the pre-trained model for text encoding in our framework. The memory buffer, denoted as $M_{t}$, retains samples from the training sets of prior tasks. This auxiliary setup serves as a valuable mechanism for enhancing task-specific learning and contributes to the robustness and adaptability of the model across different tasks.
Representation Regularization
To cope with catastrophic forget during continually debiasing pre-trained language models employing continual learning, we propose a regularization approach encompassing both generic representations $g$ and bias-specific representations $s$. This regularization is strategically applied to different representations on $g$ and $s$, as preserving knowledge in distinct spaces separately serves to encourage positive transfer and mitigate forgetting. Prior to initiating the training of all modules on debiasing task $t$, we calculate the generic representations and bias-specific representations for all sentences within the training set $D_{t}$ of the current task and the memory buffer $M_{t}$ ($M_{t}\subseteq M$). Specifically, the encoder $B^{t-1}(\cdot)$ and two disentangled networks $G^{t-1}(\cdot)$ and $B^{t-1}(\cdot)$, $D_{t-1}$ trained on the preceding task $t-1$ is employed to compute the generic representation as $G^{t-1}(B^{t-1}(x))$ and the bias-specific representation as $S^{t-1}(B^{t-1}(x))$, thereby preserving generic knowledge from the prior models to alleviate forgetting. These computed generic and bias-specific representations are retained with the representations obtained at task $t$ to perform L2 norm regularization, constraining the updated degree of the generic representation space and the bias-specific representation space. Thus, during the learning for training pairs in debiasing task $t$, we impose two regularization losses respectively:
$\mathcal{L}_{g}=E_{x\in D_{t}\cup M_{t}}||G^{t-1}(B^{t-1}(x))-G^{t}(B^{t}(x))||_{2},$ (7)
$\mathcal{L}_{s}=E_{x\in D_{t}\cup M_{t}}||S^{t-1}(B^{t-1}(x))-S^{t}(B^{t}(x))||_{2},$ (8)
where $||\cdot||_{2}$ is the L2 norm regularization operation.
A4.3 Text Classification
where $y$ denotes the corresponding class label assigned to input $x$, and $f_{cls}(\cdot)$ represents the class predictor function. The symbol $\oplus$" signifies the concatenation operation applied to the two representations under consideration.
A4.4 Text Debiasing
In the $t$-th debiasing stage, the model is aimed at debiasing the sensitive attribute $A_{t}$. Specifically, for a given sentence $x$, we adopt the given debiaser $H_{t}(\cdot)$ to mitigate sensitive attribute $A_{t}$ in the text. The debiaser $H_{t}(\cdot)$ is designed to be highly flexible and extensible, allowing for modifications as needed. In our experiments, we utilize four debiasing methods as the backbones of the debiaser $H_{t}(\cdot)$, namely FGM [15], PGD [16], ATC [17, 18, 19] and CL [20, 21, 22]. We minimize the debiaser’s loss:
$\mathcal{L}_{d}=E_{(x,y)\in D_{t}\cup M_{t}}\mathcal{L}_{d}(H_{t}(x)).$ (10)
A4.5 Memory Replay
$M_{t}^{r}=\{x\in D_{r}|P(x)\textless\beta\}(M_{t}^{r}\subseteq M_{t}).$ (11)
A4.6 Overall Objective
The final loss of our dynamic debiasing model is:
A4.7 Generalizability of Our Framework
Appendix A5 Debiasing Methods
We utilize four debiasing methods in the continuous debiasing task, namely the unsupervised debiasing methods Fast Gradient Method (FGM) [15] and Projected Gradient Descent (PGD) [16], as well as the supervised debiasing method, Adversarial Training Classifiers (ATC) [17, 18, 19] and Contrastive Learning (CL) [20, 21, 22].
Fast Gradient Method (FGM): One of the most common methods for debiasing models is unsupervised adversarial learning [57]. We follow the adversarial training method of FGM [15] to add perturbation to BERT’s representation module.
Projected Gradient Descent (PGD): PGD [16] is another unsupervised adversarial learning method, which adds an iterative attack that differs from FGM. The method performs multiple iterations, taking small steps each time, and projecting the perturbation within a specified range.
Adversarial Training Classifiers (ATC): Unlike unsupervised adversarial training, supervised adversarial training [17, 18, 19] tend to train classifiers to predict sensitive attributes for adversarial learning, thereby blurring the model’s identification of sensitive attributes.
Contrastive Learning (CL): Supervised contrastive learning [20] is utilized as an intermediate pre-training approach for mitigating biases in contextual representations. The contrastive learning [21, 22] is aimed at pulling identical entailment pairs along opposite sensitive attribute directions closer.
Appendix A6 Baselines
We use seven baselines Fine-tune [23], Experience Replay (ER) [24, 11], IDBR [25], Multi-task Learning (MTL) [26] and Two-level Knowledge Distillation (TKD) [27], Representation Projection Invariance (REPINA) [28] and (Instance-wise Relation Distillation (IRD) [29] to combine debiasing methods to verify the effectiveness of various strategies and our proposed method. Given that our experimental setting involves attribute-incremental continual learning, we transform the attribute-incremental setting into a debiasing-task-incremental setting to enable the application of the task-incremental continual learning method to our dataset.
Fine-tune [23] modifies the parameters of the pre-trained language model sequentially without any augmented strategies and additional loss.
Experience Replay (ER) [24, 11] is a memory-based technique that involves retaining a limited subset of samples from prior tasks and subsequently reintroducing them during training. The method effectively tackles the issue of catastrophic forgetting, ensuring that the model retains and integrates old knowledge in the learning process.
IDBR [25] is a novel approach that incorporates information disentanglement regularization into the encoding process. This technique disentangles the encoding of task-specific information and general information. By jointly considering these distinct encodings during the classification process, IDBR offers a promising strategy for enhancing the model’s discriminative capabilities and overall performance.
Multi-task Learning (MTL) [26] requires the model trained on all tasks simultaneously, which can effectively avoid the problem of catastrophic forgetting and be considered as a static debiasing method since it has access to data from all tasks at the same time.
Two-level Knowledge Distillation (TKD) [27] is a technique designed to improve the continual text classification process by mitigating catastrophic forgetting. This method involves transferring knowledge at two distinct levels: the feature level and the output level. The TKD strategy operates within a teacher-student framework, where the teacher is the model trained on previous tasks, and the student is the current model being trained. By integrating both feature-level and output-level distillation, the method effectively slows down the forgetting process.
Representation Projection Invariance (REPINA) [28] is a novel regularization method designed to mitigate the issue of representation collapse during the fine-tuning of pre-trained language models. REPINA aims to preserve the information content of representations by discouraging undesirable changes. This approach involves applying an invariance loss on a tunable projection of the representation, allowing it to change mildly without losing its expressivity.
Instance-wise Relation Distillation (IRD) [29] is a regularization technique used in continual learning to mitigate the forgetting of previously learned knowledge when the model learns new tasks. Specifically, it measures and minimizes the discrepancy between the representations of the current model and those of the previous model. This helps maintain the stability of the learned knowledge. By doing so, the model can adapt its representations while learning new tasks without significantly deviating from the previously learned knowledge.
Appendix A7 Supplementary Experiments
A7.1 Ablation Results
Table A1: Ablation experiment results. Upon removal of the BIR module, both the performance and fairness metrics of the models deteriorated.
@@ 4 Conclusion
As shown in Table A1, we conduct ablation experiments to verify the effectiveness of BIR. It can be seen that when ablating BID, the AAB values of CLF with ATC and CLF with CL increased by 0.0083 and 0.0096, reflecting that the model is more biased. The result demonstrates that BIR can effectively alleviate the continual biases in the model. Simultaneously, BIR would affect the performance of the model in downstream tasks to a certain extent. When discarding BID, the accuracy of CLF with ATC and CLF with CL increased by 0.0188 and 0.0028.
A7.2 Performance of Different Task Sequences
Table A2: The experimental results of the CLF with CL. The columns and rows of the table are the evaluation metrics and abbreviations for each task sequence respectively. The column located between the dotted and solid lines represents the average of the six series preceding the dotted line.
Sequence Performance Fairness
aceg 81.15 77.98 1.09 3.94 2.98 4.73 3.18
acge 87.80 84.54 4.13 8.12 6.36 5.40 6.00
aecg 86.26 82.89 3.02 7.12 6.08 5.18 5.35
aegc 85.15 81.58 2.27 7.04 4.12 6.19 4.91
agce 84.35 80.07 0.39 4.74 1.66 4.18 2.74
agec 81.00 77.45 0.14 4.66 3.34 3.42 2.89
[…]
caeg 84.89 81.44 4.93 7.53 4.61 6.92 6.00
cage 84.26 80.39 3.83 4.43 0.41 2.79 2.87
ceag 84.42 80.60 5.83 6.04 4.38 5.10 5.34
cega 84.55 81.18 6.61 8.92 4.56 5.58 6.42
cgae 76.89 73.84 1.38 3.53 3.81 4.05 3.19
cgea 84.65 81.17 2.99 8.42 5.45 8.52 6.35
[…]
eacg 85.77 82.24 3.70 6.53 1.59 3.48 3.82
eagc 86.26 83.08 2.94 8.36 6.10 3.45 5.21
ecag 84.73 81.70 2.32 7.09 5.56 5.47 5.11
ecga 86.22 82.33 2.43 6.29 3.29 4.03 4.01
egac 80.31 76.69 0.35 4.28 1.56 5.41 2.90
egca 80.48 77.64 2.06 5.48 3.20 4.06 3.70
[…]
gace 83.78 79.61 1.64 5.15 2.09 4.49 3.34
gaec 85.08 80.90 2.69 5.45 2.18 4.09 3.60
gcae 85.61 81.72 2.86 6.45 4.21 4.04 4.39
gcea 86.21 81.79 4.51 8.83 3.43 5.49 5.56
geac 79.62 75.80 3.74 4.84 1.80 4.25 3.66
geca 85.08 80.76 1.22 5.83 3.82 5.78 4.16
[…]
Avg 83.94 80.31 2.79 6.21 3.61 4.84 4.36
We further explore the impact of the debiasing task order on the performance and fairness of the model. The experiment results of 24 task sequences for the CLF with CL model are shown in Table A2. When the debiasing order is “agce”, the average bias of the four sensitive attributes in the model is the smallest, with an ABB value of 0.0274, while concerning debiasing order “cgea”, the ABB value reaches 0.0635. We further calculate the average ABB values when the initial debiasing attribute in the debiased sequence is “age”, “country”, “ethnicity”, and “gender”. The results reveal that the selection of the initial debiasing attribute will affect the fairness of the model. When first mitigating the biases related to the sensitive attribute “country”, the model retains more biases. It can also be seen that in this case, the average performance of the model in downstream tasks is the worst. It is possible that the features associated with the country attribute have a great impact on the model’s ability to recognize hate speech. Therefore, when the model initially debiases the country attribute and subsequently forgets its debiasing ability, it amplifies the bias of this attribute.
A7.3 Experiments on Other Pre-trained Models
Table A3: Results of Deberta-v3-base and RoBERTa-base. The columns and rows of the table are the evaluation metrics and the names of each method respectively. Experimental results on different backbones show that our method has obvious continuous debiasing capabilities and is better than existing methods.

### v2

@@ V-B Debiasing Performance Forgetting of Continuous Debias- ing
As shown in Table V, we analyze the results to explore the impact of persistent bias on the fairness of models and whether the debiasing performance of the model suffers from the catastrophic forgetting problem in the continuous debiasing task. Compared MTL-based methods representing static debiasing methods with Fine-tune-based methods indicating continuous debiasing methods, the results reveal that the AAB of the model under continuous debiasing tasks is consistently higher than that of static debiasing setting, demonstrating stronger bias. This indicates that the continuous debiasing performance of models may suffer from catastrophic forgetting with ever-changing biases. The HateDebias dataset closely mirrors real-world bias occurrences, validating that static methods are inadequate under continual biases. Thus, it is significant to develop adaptive debiasing methods that enhance the overall performance and fairness of models in practical applications. This finding underscores the rationale and necessity of the HateDebias dataset proposed in this paper. It closely resembles the occurrence of bias in real-world scenarios, simulating multiple, continuously changing biases, and provides a real environment closer to practical applications.
V-C Performance of Different Methods on HateDebias
@@
As detailed in Table V, our experimental findings indicate that our approach yields substantial improvements for the fairness of models in continual debiasing tasks. Static debiasing methods (MTL), which can be considered as an upper-bound followed by [46, 50], show significant debiasing effects when dealing with multiple sensitive attributes, while fine-tuning-based continuous debiasing methods have limited debiasing capabilities. The models with fine-tuning-based methods are less stable in debiasing and less effective in addressing continuous bias, with larger AAB values and the variances. Additionally, the continuous debiasing effect using the ER, REPINA and IRD methods is worse, indicating an inability to effectively reduce bias. These methods are ineffective because the HateDebias dataset closely mirrors real-world bias occurrences, simulating multiple and continuously changing biases. Experimental results show that existing methods perform poorly in these complex, dynamic environments. Finally, the experimental results show that our proposed framework significantly reduces bias in continuous debiasing tasks. Especially with supervised debiasing strategies, the model’s bias is notably reduced at the cost of slight performance trade-offs in downstream tasks. Compared to methods like REPINA, IRD, and TKD, the CLF method demonstrates better debiasing effects and model performance across various debiasing strategies. Our results demonstrate that CLF achieves the best trade-offs as measured by DTO, outperforming other methodologies included in the study. Although CLF shows lower raw performance in some metrics, it significantly improves fairness. Experimental results show the proposed method CLF excels in handling multiple and continuously changing biases in complex real-world scenarios, effectively debiasing while maintaining high performance and fairness.
V-D Performance of Debiasing Preservation Capability
$\small BC=\frac{1}{n-1}\sum_{i=1}^{n-1}FPR_{i}^{n}-FPR_{i}^{i},$ (9)
where $FPR_{i}^{i}$ is the false positive rate of the $i$-th bias in the $i$-th stage. Table VI presents the results of the debiasing preservation capability. These results demonstrate that our proposed CLF method significantly outperforms others in preserving debiasing effects over time. Specifically, the CLF method achieves a BC value of -1.60, indicating substantial bias reduction and showcasing its effectiveness in mitigating bias. In comparison, the REPINA method shows no improvement, while methods like TKD and IDBR exhibit moderate improvements. The poor performance of the comparative methods indirectly reflects the reasonableness of our dataset in terms of capturing the diversity and variability of real-world biases.
TABLE VI: The results of debiasing preservation capability.
@@ V-D Performance of Debiasing Preservation Capability
TABLE VII: Experimental results of different sequence lengths.
@@
TABLE VIII: Ablation experiment results. Upon removal of the BIR module, both the performance and fairness metrics of the models deteriorated.
@@
TABLE IX: Results of Deberta-v3-base and RoBERTa-base. The columns and rows of the table are the evaluation metrics and the names of each method respectively. Experimental results on different backbones show that our method has obvious continuous debiasing capabilities and is better than existing methods.
@@
V-E Effects of Different Length Settings on HateDebias
In practical applications, biases come in various types and forms. To gradually make our dataset approximate real-world scenarios, we experiment with different sequence lengths to evaluate the debiasing effects under different length settings. We compare continuous debiasing tasks with sequence lengths of 2, 3 and 4 to analyze how the diversity and variability of the dataset affect model performance and fairness. As shown in Table  VII, we choose CL as the debiasing method and the results reveal several key insights: (1) As the sequence length increases, the performance metrics generally show a decreasing trend. This suggests that longer sequences may introduce more complexity, making it challenging for models to maintain high performance. (2) The AAB values decrease with longer sequence lengths, indicating improved fairness. This trend demonstrates that datasets with longer sequences better capture the diversity and variability of real-world biases, helping models to generalize and debias more effectively. (3) Our proposed method CLF exhibits the best continuous debiasing capability across different sequence lengths, indicating that it can effectively cover datasets in various real-world scenarios.
V-F Ablation Results
We discard the regularization losses $\mathcal{L}_{g}$ and $\mathcal{L}_{s}$ to conduct ablation experiments to verify the effectiveness of BIR, as shown in Table VIII. It can be seen that when ablating BID, the AAB values of CLF with ATC and CLF with CL increased by 0.0083 and 0.0096, reflecting that the model is more biased. The result demonstrates that BIR can effectively alleviate the continual biases in the model. Simultaneously, BIR would affect the performance of the model in downstream tasks to a certain extent. When discarding BID, the accuracy of CLF with ATC and CLF with CL increased by 0.0188 and 0.0028.
V-G Experiments on Other Pre-trained Models